Let V be the vector space of all n x n
matrices over R. What is the dimension of
V over R ? Further, let
Wn={Anxn belongs to V I Anxn is upper triangular}.
Check whether or not W is a subspace of V.
1
Expert's answer
2020-12-15T03:49:51-0500

Let "V" be the vector space of all "n \times n" matrices over "\mathbb R".  For "i,j\in\{1,2,...,n\}" denote by "E_{ij}" the matrix with a unigue 1 at the intersection of the row "i" and the column "j", and with the rest elements equal to 0: "e_{ij}=1" and "e_{st}=0" for all "s\ne i" and "t\ne j". Then "(E_{ij}\ |\ i,j\in\{1,..., n\})" is a basis of the vector space "V". Indeed, if "\sum\limits_{i=1}^n\sum\limits_{j=1}^na_{ij}E_{ij}=O" where the matrix "O" contains of only zeros, then "a_{ij}=0" for all "i,j\in\{1,2,...,n\}". Consequently, the matrices "E_{ij}" are linear independent. Also each matrix "M" with entries "m_{ij}" can be represented as "M=\sum\limits_{i=1}^n\sum\limits_{j=1}^nm_{ij}E_{ij}", and therefore, "(E_{ij}\ |\ i,j\in\{1,..., n\})" indeed a basis of "V". It follows that "\dim(V)=n^2."



Further, let "W_n=\{A_{n\times n} \in V \ |\ A_{n\times n}\text{ is upper triangular}\}." Let  "{\displaystyle A_{n\times n}, B_{n\times n}}" be elements of "W_n"  and "{\displaystyle \alpha ,\beta }"  be elements of "\mathbb R". Since the elements under the main diagonal in matrices  "{\displaystyle A_{n\times n}, B_{n\times n}}" are zeros, it follows that the elements under the main diagonal in the matrix "{\displaystyle\alpha A_{n\times n}+\beta B_{n\times n}}" is zeros as well. Therefore,  "{\displaystyle\alpha A_{n\times n}+\beta B_{n\times n}}"  is in "W_n", and we conclude that "W_n" is a subspace of "V".